OpenStudy (anonymous):

If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4).

OpenStudy (anonymous):

please help me?

Parth (parthkohli):

When \(a\) varies directly with \(b\), then \(a = cb\) where \(c\) is a constant. You understand this, right?

OpenStudy (anonymous):

yes, but I am in quadratic variation atm, so I don't klnow if it wants direct or quad

Parth (parthkohli):

Don't worry about it. It's simple. And yes, the earlier points come in fast.

OpenStudy (anonymous):

what should I start with?

Parth (parthkohli):

f(x) varies directly with x^2. So,\[f(x) = c \times x^2\]Let's do a job: find the \(c\) with the information we're given. It is given that \(f(5) = 75\). But \(f(5)\) is also the number you get when you plug in 5 for x in the expression \(c \times x^2\).

OpenStudy (anonymous):

mhm...

Parth (parthkohli):

So you get \(f(5)=c \times (5)^2=25c\). But \(f(5) = 75\) as given. Therefore,\[25c =75 \Rightarrow c = 3\]

Parth (parthkohli):

Since we're doing direct relation, the constant is always the same. So \(c=3\) in this direct relation. Thus, here's our direct relation:\[f(x)=3\times x^2\]

Parth (parthkohli):

You can now find \(f(4)\)

OpenStudy (anonymous):

brb urinating... sorry about that.

Parth (parthkohli):

...

OpenStudy (anonymous):

back

Parth (parthkohli):

I've explained everything to you. You're left with\[f(x) = 3x^2\]

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

48?

Parth (parthkohli):

You have to find f(4), so plug 4 into that.